OpenStudy (anonymous):

What is the solution for 2x^2-3x+2=0 Help me figure out how to solve ones like it too?

OpenStudy (campbell_st):

well you will need the general quadratic formula as the solutions are complex numbers.

OpenStudy (anonymous):

Okay, I will look up that formula?

OpenStudy (studygurl14):

Quadratic formula: Where you have this quadratic: \(\large ax^2+bx+c=0\) \(\large x=\Large\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

Okay, I understand that much. But as my b is already negative, do I leave it that way? Or make it -(-3) thus causing a positive?

OpenStudy (studygurl14):

If you have \(\large 2x^2-3x+2=0\) What does a, b, and c equal?

OpenStudy (studygurl14):

the second way. make it postive

OpenStudy (anonymous):

Okay. so x= 3 +- the sqrt of -3^2-4(2)(2)/4

OpenStudy (studygurl14):

so when you plug in you get \(\large x=\Large\frac{-(-3)\pm\sqrt{(-3)^2-4(2)(2)}}{2(2)}\)

OpenStudy (anonymous):

Okay, cool. I haven't screwed up yet.

OpenStudy (studygurl14):

lol...just simplify

OpenStudy (anonymous):

So 3+- sqrt of 9-16/4?

OpenStudy (studygurl14):

correct

OpenStudy (studygurl14):

\(\large x=\Large\frac{3\pm\sqrt{9-16}}{4}\)

OpenStudy (anonymous):

I'll just solve it from here and you tell me if I've done it correctly. 9-16=-7 so 3+- sqrt of -7/4?

OpenStudy (studygurl14):

correct

OpenStudy (studygurl14):

that's it. unless you need to submit your answer to the nearest tenth or something like that

OpenStudy (studygurl14):

oh, wait. not it. sorry

OpenStudy (studygurl14):

you have to take out the imaginary number

OpenStudy (studygurl14):

\(\large x=\Large\frac{3\pm\sqrt{-7}}{4}\) do you know about \(\large i\) ?

OpenStudy (anonymous):

Well, it's either 3+-i sqrt 7/4 OR 3+-5i/4

OpenStudy (anonymous):

I do know that i= sqrt of negative one.

OpenStudy (anonymous):

Which is exactly what I was going to ask you lol, how do I remove the imaginary number?

OpenStudy (studygurl14):

right

OpenStudy (studygurl14):

\(\large x=\Large\frac{3\pm\sqrt{7\times-1}}{4}\rightarrow\frac{3\pm\ i\sqrt{7}}{4}\rightarrow\)

OpenStudy (studygurl14):

ignore the last arrow on teh end

OpenStudy (anonymous):

Okay, so you turn the negative into an i?

OpenStudy (anonymous):

Ohhh! You multiply it out! Duhh! Got it. Thank you so much!

OpenStudy (studygurl14):

:)

OpenStudy (studygurl14):

anytime, my friend \(\large\color{blue}{:)}\)